OpenStudy (anonymous):

Tell what you can do to the first inequality in order to get the second. 1.) 5z>-25;z>-5 2.) -c/4>3;C<-12 Determine whether each statement is always, sometimes, or never true. justify your answer. 3.) If x>3 and y<1, then xy>0 4.) If x<0 and y<0, then xy >0 Please help me I'm stuck

OpenStudy (anonymous):

so what do you need help with>

OpenStudy (anonymous):

Can you like the whole complete sentence exactly what you see on your question?

OpenStudy (anonymous):

type the whole

OpenStudy (anonymous):

Thats the whole thing

OpenStudy (anonymous):

> are these supposed to be arrows or less than arrows?

OpenStudy (anonymous):

Less then symbols

OpenStudy (anonymous):

With this equation I would first gather the z-terms on the left of the equation to get 3z + 2z - 5 = 25 - 5z, which is 5z - 5 = 25 - 5z To eliminate the z-term from the right side of the equation, add 5z to both sides. This maintains the balance and you get 5z - 5 + 5z = 25 - 5z + 5z, or 10z - 5 = 25 Now add 5 to both sides, again maintaining the balance to get 10z - 5 + 5 = 25 + 5, or 10z = 30. Finally divide both sides by 10 to get

OpenStudy (anonymous):

Thanks! Can you help me with the other ones?

OpenStudy (anonymous):

@ImakeTyRaNiTe